Consider two individuals, Bill and Steve. Both can make cars and computers. Their production levels are outlined below:

In one hour of time Bill can make: 7 Cars -or- 35 Computers.
In one hour of time Steve can make: 4 Cars -or- 12 Computers.

1. What is Bill's opportunity cost of making one (1) Car?
What is Steve's opportunity cost of making one (1) Car?

2. What is Bill's opportunity cost of making one (1) Computer?
What is Steve's opportunity cost of making one (1) Computer?

3. a. Who has the absolute advantage in Car production? Explain.
b. Who has the absolute advantage in Computer production? Explain.

4 a. Who has comparative advantage in Car production? Explain.
b. Who has comparative advantage in Computer production? Explain.

5. In the event the two decide to specialize and trade away what they are better in making, who will produce what? What will the terms of trade be? Who is going to benefit and by how much? Please be specific.
Answer #1

1) BIlls , opportunity of 7 cars = 35 computers

Bills Opportunity cost of making one car = what he is sacrifieng / what he is gainning

= 35 / 7

= 5

steve can make 4 cars or 12 computer in 1 hour

Opportunity cost of making one car = what he is sacrifieng / what he is gainning

= 12 / 4

  Opportunity cost of making one car = 3

2)bills 35 computer = 7 cars

bobs Opportunity cost of making one computer = what he is sacrifieng / what he is gainning

= 7 / 35

= 1 /5

steve , 12 compuerters = 4 cars

steve's Opportunity cost of making one computer = what he is sacrifieng / what he is gainning

= 4 / 12

= 1 /3

3 ) a ) to know the absolute advantage we compare productivity of 2 individuals

in 1 hour bill produced 7 cars and Steve produces 4 cars  so Bill has absolute advantage in Car production as productivity of making cars of Bill is high .

B) in 1 hour bill produced 35 computers and Steve can produces 12 compurets so here also Bill has absolute advantage in Car production.as productivity of making computers of Bill is high .

4)a) to know the comprative advantage we should first know the opportunity cost

as we already solved opportunity cost in question 1

Steve has comprative advantage in car production as steve's opportunity  making one car = 3 while Bills Opportunity cost of making one car = 5

b) Bill has comprative advantage in computer  production

as bills opportunity cost of making one computer =1/5 , while steve's Opportunity cost of making one computer = 1/3

5) if they want to specialise then they will both choose the quantity in which they are compratively advantage

Bills has comprative advantage in computer production so he will produce = 35 computers

Steve has comprative advantage in car production so he will produce = 4 cars

Bills Opportunity cost of making one car = 5 , so he will be willing to pay maximum of 5  computers to Steve for 1 car

steve's Opportunity cost of making one computer = 1 / 3 so he will be willing to exchange 1 car for minimum of 3 computers.